A vendor buys orange at 26per dozen and sells them at 5for 13 find his gain percent

Grade:7

1 Answers

RUSHIKESH RAMESHCHANDRA BENKE
2 years ago
One dozen means 12.
Vendor buys 12 oranges for 26.
That means 1orange for (26/12)
And sell 5 oranges for 13.That means 1 orange for (13/5)
Now Gain formula is=(Sell Prize-Cost Prize)/Cost prize
Now put the values in formula
Gain=((\frac{13}{5}-\frac{26}{12})/\frac{26}{12})*100
        =\frac{26}{130}*100
        =20 percent
A vendor gain 20 percent